OpenStudy (anonymous):

steps to calculate Integral of dx/(9+4x^2)

OpenStudy (anonymous):

what ways of integration have you learned?

OpenStudy (uri):

@toby-san You're in which grade?

OpenStudy (raden):

i mean, let x = (3/2)tanθ :)

OpenStudy (raden):

x = (3/2)tanθ dx = (3/4) sec^2 (θ) dθ see ur integral becomes int (3/4) sec^2 (θ)/3secθ dθ = int 1/4 secθ dθ

OpenStudy (raden):

now, can evaluate that integral ?

OpenStudy (raden):

* can u

OpenStudy (raden):

hint : multiple by (secθ+tanθ)/(secθ + tanθ) then integrate by u-subt

OpenStudy (raden):

aha, i thought it is sqrt(9+4x^2) . sorry :) try again : x = (3/2)tanθ dx = (3/4) sec^2 (θ) dθ int (3/4) sec^2 (θ)/9sec^2 θ dθ = int 1/12 dθ = θ/12 + c

OpenStudy (raden):

damn, mistake again x = (3/2)tanθ dx = (3/2) sec^2 (θ) dθ int (3/2) sec^2 (θ)/9sec^2 θ dθ = int 1/6 dθ = θ/6 + c because x = (3/2)tanθ, so tanθ = 2x/3 θ = tan^-1 (2x/3) so, it is 1/6 tan^-1 (2x/3) + c